OpenStudy (anonymous):

write an equation of a line that passes through (-6,2) with slope 2/3

OpenStudy (anonymous):

Use the point-slope form of an equation of a line: y-y1= m(x-x1)

OpenStudy (anonymous):

wait how would i put them in here?

OpenStudy (anonymous):

oh sry, i thought ur slope was a coordinate poitn...my bad sry

OpenStudy (anonymous):

i get messed up with the y and x things

OpenStudy (anonymous):

-6 is ur x 2 is ur y

OpenStudy (anonymous):

so then i have 2-y^1=m(-6-x^1)

OpenStudy (anonymous):

Slope-intercept form is like this: y = mx + b where m is the slope, and b is the y intercept. So they give you the slope of 2/3 which means m = 2/3, right? They also give you point (-6,2) which means they have given you an x-value and a y-value for this line: x = -6 and y = 2. Good so far?

OpenStudy (anonymous):

im not sure this is the answers i have A)y=2/3x+8 B)y=2/3x+6 C)y=2/3x-2 D)y=2/3-4

OpenStudy (anonymous):

what i'm asking, is if you understand what i've written so far:)

OpenStudy (anonymous):

o yeah i understand

OpenStudy (anonymous):

sorry got confushed

OpenStudy (anonymous):

what are you confused on?

OpenStudy (anonymous):

i dont know now

OpenStudy (anonymous):

I'll start over then:) (This is really important to understand because it's the basics)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So the choices you are given are all in slope intercept form. (y = mx + b) Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay:) And you are given the slope = 2/3 and the point = (-6,2) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the equation y = mx + b is what we'll plug that info into okay? SO if m = the slope and b is the y intercept (the y part of the point they gave you) can you plug it in?

OpenStudy (anonymous):

2=-6x+b? is this how i would plug them in?

OpenStudy (anonymous):

um no. the only thing you're pluggin in is the slope where the m is and the y intercept where the b is :)

OpenStudy (anonymous):

can you try again?

OpenStudy (anonymous):

y=2/3x+-6?

OpenStudy (anonymous):

wait. you are VERY close. You accidentally put the x- intercept in there instead of the y - intercept. Try once more :D

OpenStudy (anonymous):

y=2/3x+2

OpenStudy (anonymous):

YAY! There you go:D

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

would the +2 turn to -2?

OpenStudy (anonymous):

Only if the 2 was a negative.

OpenStudy (anonymous):

o, i dont have this answer to choice from

OpenStudy (anonymous):

hmm. that's odd. are you sure you wrote this part right "(-6,2)" ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i have y=2/3x+8 y=2/3x+6 y=2/3x-2 y=2/3x-4

OpenStudy (anonymous):

did you post the question right though?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the slope supposed to be 2/3?? or -2/3?

OpenStudy (anonymous):

write an equation of a line that passes through (-6,2) with slope 2/3.

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

okay then. we can try another equation. y-y1= m(x-x1) with this equation, you plug in the 2 for the y1, and the -6 for the x1. can you do that for me?

OpenStudy (anonymous):

If you plug it in, it should look like this: y-2= m(x+6) then you plug the slope (2/3) in with the m, and get y-2= 2/3(x+6) from here, you solve. do you understand so far?

OpenStudy (anonymous):

i thought the 6 would be -

OpenStudy (anonymous):

yay! it would (i was just seeing if you really cared about understanding it) So it's y-2= 2/3(x-6) now if we further solve it we get y - 2 = 2/3x - 4 we add 2 to both sides and the 4 turns into a 2 and we get what?

OpenStudy (anonymous):

y=2/3x+6?

OpenStudy (anonymous):

no the answer is y=2/3x-2

OpenStudy (anonymous):

yay:) there you go :D sorry about the confusion about the last one.

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

(you're welcome i guess then?)